A student takes a 20-question, multiple choice exam with five choices for each question and guesses on each question. Assuming for each question only one of the five choices is correct, what is the mean number of questions the student could get correct by guessing?
Answer by stanbon(48568)

A student takes a 20-question, multiple choice exam with five choices for each question and guesses on each question. Assuming for each question only one of the five choices is correct, what is the mean number of questions the student could get correct by guessing?
--------------------------------------
Binomial Problem:
n = 20 ; p = 1/5
----
mean = np = (1/5)20 = 4
---------------------------------
Cheers,
Stan H.